Lulabalula:

Today, Reza ate 20 grapes, which was 400% as many grapes as he ate yesterday. Explain how to use equivalent ratios to find the number of grapes Reza ate yesterday.

Janiah:

why just why

Lulabalula:

idk

jhonyy9:

using the rule of 3 simple you can write yesterday ate x what was 100% today ate 20 what was 400% and now cross multiplie and calcule the x value x = 20*100 /400 = 20/4 = ? hope helped understandably easy